Two cars A and B are at a positions 100 m and 200 m from the origin at time t=0. They start simultaneously with velocities 10m/s & 5m/s respectively. Determine the position of time at which they will overtake one another. Assume that they are moving in the same direction. 

Grade:11

1 Answers

Arun
5 years ago
Let S be the distance from the pont car A start to the distance where they both meet and t be the time taken to by both car to meet

Car A:
 
S=10t

Car B:

S-100 = 5t

dividing both equation we get

2S-200=S
S=200 

ie... they both meet at distance of 200m from the start point of A ... which implies the distance from origin is 200+100 = 300m

time taken =>

S=10t
t=20seconds

they take 20 seconds to meet